OpenStudy (anonymous):

Help!!!??? The US Maritime Administration estimated that the cost per ton of building an oil tanker could be represented by the model y=113,000/x+235, where y is the cost in the dollars per ton and x is the tons(in thousands). What size of oil tanker (in thousands of tons) can be uilts for $200 per ton?

hero (hero):

Hint: y = 200

OpenStudy (anonymous):

Do I just divide 113,000?

hero (hero):

You have to solve for x

hero (hero):

200 = 113,000/x + 235

OpenStudy (anonymous):

then do I multiply it?

OpenStudy (anonymous):

\[200=\frac{ 113000 }{ x+235 }\]

hero (hero):

So, yes multiply both sides by x + 235

hero (hero):

Let me know what you get afterwards.

OpenStudy (anonymous):

200x+47000=113000+26555000

OpenStudy (anonymous):

But solving it that way just got me x=-235

hero (hero):

That can't be right

OpenStudy (anonymous):

I know...now I'm just confused

hero (hero):

That's not the right value for x

hero (hero):

x is positive anyway

OpenStudy (anonymous):

Yes, I know that, but now I'm not sure how to get x

hero (hero):

Yeah, you messed up while solving for x somehow.

OpenStudy (anonymous):

Well I told you what I got when I multiplied both sides....was that wrong?

hero (hero):

(x + 235)/(x+235) should have cancelled on the right side.

hero (hero):

Leaving you with just 200x+47000=113000 now finish solving for x

OpenStudy (anonymous):

Ohh, I see what I did wrong. Thank you!

OpenStudy (anonymous):

There is also one equation that I just can't figure out.